On June 30, 2012, Pharoah Company issued 12% bonds with a par value of $860,000 due in 20 years. They were issued at 98 and were callable at 104 at any date after June 30, 2020. Because of lower interest rates and a significant change in the company’s credit rating, it was decided to call the entire issue on June 30, 2021, and to issue new bonds. New 8% bonds were sold in the amount of $980,000 at 103; they mature in 20 years. Pharoah Company uses straight-line amortization. Interest payment dates are December 31 and June 30.

(a) Prepare journal entries to record the redemption of the old issue and the sale of the new issue on June 30, 2021.
(b)

Prepare the entry required on December 31, 2021, to record the payment of the first 6 months’ interest and the amortization of premium on the bonds.

Account Titles and Explanation Debit Credit No. Date (a) June 30, 2021 Bonds Payable M0.000 Cash 894 400 (To record the redemplease provide calculations. thank you.
Answer #1
Discount on issue of bonds 17200 =860000*(1-0.98)
Discount amortized for 9 years 7740 =17200*9/20
Unamortized discount 9460 =17200-7740
Cash paid for redemption 894400 =860000*1.04
Less: Carrying value of bonds 850540 =860000-9460
Loss on redemption 43860
Date Account Titles and Explanation Debit Credit
June 30,2021 Bonds payable 860000
Loss on redemption of bonds 43860
      Discount on Bonds payable 9460
      Cash 894400
(To record the redemption of the old issue)
June 30,2021 Cash 1009400 =980000*1.03
      Bonds payable 980000
      Premium on Bonds payable 29400
(To record the sale of the new issue)
b
December 31,2021 Interest expense 38465
Premium on Bonds payable 735 =29400/20*6/12
      Cash 39200 =980000*8%*6/12